OpenStudy (anonymous):

Allie has a birthday present for her brother. The length of the box that the present comes in is two times longer than the width of the box. The height of the box is 5 inches and its volume is 160 cubic inches. What is the length and width of the box?

OpenStudy (anonymous):

Volume=L*W*H You know the volume and height. The length can be expressed as 2 times the width or 2w. And the width as just w. 160=2w*w*5 Solve for w

OpenStudy (anonymous):

How do i solve for w?

OpenStudy (anonymous):

First multiply the right side and you get: 160=10w^2 Divide both sides by 10 16=w^2 Take the square root of both sides \[\sqrt{w^2}=\sqrt{16}\] w=4

OpenStudy (anonymous):

So since the length is 2 times the width, the length is 2*4

OpenStudy (anonymous):

width=4 length=8